OpenStudy (anonymous):

The average life of a dog is 4 times as great as the averages life span of a mouse. If the average life span of a dog is 12 years, write and solve an equation to find the average life span of a mouse.

OpenStudy (anonymous):

Let the life span of a mouse be x. 4x = 12 x = 3 years

OpenStudy (anonymous):

d=4m m=d/4 m=12/4 m=3

OpenStudy (king):

x=4y 12=4y y=3

OpenStudy (anonymous):

Thanks! guys...

OpenStudy (anonymous):

yw